OpenStudy (anonymous):

factor: x^2+y-xy-x ??

OpenStudy (anonymous):

x^2+y-xy-x =x^2-x+y-xy =x(x-1)+y(1-x) =x(x-1)-y(x-1) =(x-y)(x-1)